# LASIL: Learner-Aware Supervised Imitation Learning For Long-term Microscopic Traffic Simulation

### v1

@@ 1 Introduction
However, creating a realistic simulator that can simultaneously replicate the microscopic response of human drivers in various traffic conditions and long-term macroscopic traffic statistics is challenging. In recent years, there have been significant efforts to develop realistic traffic simulators that aim to accurately model human driving behaviors. Traditional traffic simulators, such as SUMO [23], AIMSUN [2], and MITSIM [39], typically rely on heuristic car-following models like the Intelligent Driver Model (IDM) [38]. However, despite carefully calibrating parameters, these simplified, rule-based models often fail to deliver accurate simulations due to the complexity of real-world traffic environments [11, 27]. Factors such as the road structure, neighboring vehicles, and even driver psychology can influence the decisions of human drivers, making it challenging to achieve accurate simulations.
In pursuit of realistic traffic simulation, researchers have turned to neural networks to represent the driving model through imitation learning (IL) from human demonstrations. Most traffic simulation approaches [31, 4, 37] leverage behavior cloning (BC) [29] to learn a driving policy by minimizing the disparity between the model output and the human demonstrations in training data. However, the BC method is hindered by covariate shift [34], where the state induced by the learner’s policy progressively diverges from the expert’s distribution. Existing BC-based simulators have succeeded in short-term (no more than 20 seconds) simulation applications like autonomous driving tests but often fail to generate stable long-term simulations.
To address covariate shift, existing methods such as DAgger [34] and DART [24] incorporate supervisor corrections at the learner’s or perturbed expert’s states. However, collecting human supervision data can be challenging due to intensive labor and judgment errors. Recent traffic simulators [36, 41] propose using inverse reinforcement learning (IRL). These methods, such as generative adversarial imitation learning (GAIL) [16] and adversarial inverse reinforcement learning (AIRL) [12], learn a reward function using a discriminator neural network within Generative Adversarial Networks (GANs) [13]. The policy network is trained to maximize the learned reward through online reinforcement learning (RL), enabling agents to handle out-of-distribution states. However, directly applying GAIL to traffic simulation can be problematic [5]. The dynamic nature of the environment in the multi-agent system can lead to noises during policy learning, resulting in highly biased estimated gradients. Furthermore, training the discriminator in GAIL is challenging due to the instability and sensitivity of hyperparameters during min-max optimization.
@@ 2.1 Imitation learning
Due to the challenges in obtaining expert supervision, recent IRL-based methods utilize feedback from a neural network-based discriminator to handle covariate shift. Typically, these methods involve an iterative process alternating between reward estimation and reinforcement learning. Earlier IRL-based methods [32, 42, 14] require frequent dynamic programming processes, while recent adversarial IL approaches integrate reward function learning with policy learning using a GAN formulation. However, both GAN and RL training processes are known to be unstable, sensitive to hyperparameters, and have poor sample efficiency [8]. Moreover, the discriminator can easily exploit insignificant differences between expert and policy samples, leading to undesirable performance [40]. In contrast, our method avoids the min-max optimization problem and the sample-inefficient RL process, requiring minimal fine-tuning. Instead, we utilize the real future trajectory as the target state to learn corrective behavior.
@@ 2.2 Imitation Learning-based Traffic simulation
Recent advancements in traffic simulators have focused on enhancing realism by leveraging imitation learning (IL) from human driving demonstrations. These simulators extend traditional BC and IRL methods to tackle the challenging multi-agent IL problem.
@@ 5.2 Metrics
We evaluate the realism of our simulator by measuring the similarity between the simulation result and real data. During evaluation, we assume that each vehicle enters the simulator at its first recorded time and position, and is then controlled by our simulator to complete its recorded route. When an agent reaches its final recorded position, it is removed from the simulator.
Short-term microscopic: Following prior works [36, 5], we conduct a short-term microscopic evaluation by simulating for 20 seconds from a random time step in the test dataset. We first measure the similarity between the simulated and real data using position and velocity RMSE metrics, which are calculated by:
@@
5.3 Performance
We compare our method against state-of-the-art baselines:
BC [29]: we learn our model directly by the BC method.
@@ 5.4 Ablation Study
We conducted a series of ablation experiments to assess the individual contributions of crucial components in our approach. The results, presented in Tab. 1 and Tab. 2, highlight the impact of each component on the performance of our traffic simulation model.
@@
On-road projection: Our ablation study on the on-road projection module aims to show its impact on reducing the off-road rate. The results demonstrate that the on-road projection module leads to a notable decrease in the off-road rate and moderate improvements in other performance metrics.
LQR: Removing the LQR module allowed us to evaluate its effectiveness. While the LQR led to a higher short-term off-road rate due to more constrained movement (driving off-road due to inertia), its removal deteriorated other short-term and long-term metrics. The reason is that the LQR can smooth the trajectory, thus leading to more realistic simulation.
@@ 5.5 Qualitative Result
In Fig. 2, we present the mean road density and speed for real-world data, SUMO simulation, and our proposed method over all time steps. The results demonstrate that our proposed method excels at replicating long-term macroscopic traffic patterns, surpassing the capabilities of the SUMO simulator. This achievement can be attributed to our model’s enhanced ability to replicate the typical microscopic driving behaviors over a long period.
@@ 7 Model details
The hyper-parameters of our model architecture are listed in Tab. 3. All models are trained using the Adam optimizer with a learning rate of 0.0003, and batch size 4 with 8 Ge-force 3090. The VAE are trained simultaneously with the policy network using the online simulated learned data and offline expert data.

### v3

@@ 1 Introduction
However, generating realistic and accurate simulations that can simultaneously replicate the microscopic response of human drivers in various traffic conditions and long-term macroscopic traffic statistics is challenging. In recent years, significant efforts have been invested in developing realistic traffic simulators with the goal to accurately model human driving behaviors. Traditional traffic simulators, such as SUMO [24], AIMSUN [2], and MITSIM [44], typically rely on heuristic car-following models like the Intelligent Driver Model (IDM) [42]. However, despite carefully calibrating parameters, the rule-based models often fail to deliver accurate simulations due to the complexity of real-world traffic environments [12, 30]. Factors such as the road structure, neighboring vehicles, and even driver psychology can influence the decisions of human drivers, making it challenging to achieve accurate simulations [43, 27, 7].
In pursuit of realistic traffic simulations, researchers have turned to neural networks to represent the driving model through imitation learning (IL) from human demonstrations. Most traffic simulation approaches [34, 4, 41] leverage behavior cloning (BC) [32] to learn a driving policy by minimizing the disparity between the model output and the human demonstrations in training data. However, BC is hindered by covariate shift [37], where the state induced by the learner’s policy progressively diverges from the expert’s distribution. Existing BC-based simulators have succeeded in short-term (less than 20 seconds) simulation applications like autonomous driving tests but often fail to generate stable long-term simulations.
To address covariate shift, existing methods such as DAgger [37], DART [25], and ADAPS [28] incorporate supervisor (humans or principled simulations) corrections at the learner’s or perturbed expert’s states. However, human supervision can be problematic due to intensive labor and judgment errors, and principled simulations may not account for heterogeneous driving behaviors. Recent traffic simulators [40, 46] propose using inverse reinforcement learning (IRL). These methods, such as generative adversarial imitation learning (GAIL) [17] and adversarial inverse reinforcement learning (AIRL) [13], learn a reward function using a discriminator neural network within Generative Adversarial Networks (GANs) [14]. The policy network is trained to maximize the learned reward through online reinforcement learning (RL), enabling agents to handle out-of-distribution states. However, directly applying GAIL to traffic simulation can be problematic [5]. The dynamic nature of the environment in the multi-agent system can lead to noises during policy learning, resulting in highly biased estimated gradients. Furthermore, training the discriminator in GAIL is challenging due to the instability and sensitivity of hyperparameters during min-max optimization.
@@ 2.1 Imitation learning
Due to the challenges in obtaining expert supervision, recent IRL-based methods utilize feedback from a neural network-based discriminator to handle covariate shift. Typically, these methods involve an iterative process alternating between reward estimation and reinforcement learning. Earlier IRL-based methods [35, 47, 39, 15] require frequent dynamic programming processes, while recent adversarial IL approaches integrate reward function learning with policy learning using a GAN formulation. However, both GAN and RL training processes are known to be unstable, sensitive to hyperparameters, and have poor sample efficiency [9]. Moreover, the discriminator can easily exploit insignificant differences between expert and policy samples, leading to undesirable performance [45]. In contrast, our method avoids the min-max optimization problem and the sample-inefficient RL process, requiring minimal fine-tuning. Instead, we utilize the real future trajectory as the target state to learn corrective behavior.
@@ 2.2 Imitation Learning-based Traffic simulation
Recent traffic simulators focus on enhancing realism by leveraging imitation learning (IL) from human driving demonstrations, which extend traditional BC and IRL methods to tackle the challenging multi-agent IL problem.
@@ 5.2 Metrics
We evaluate the realism of our simulator by measuring the similarity between the simulation result and real data. During evaluation, each vehicle enters the simulator at its first recorded time and position, and is controlled by our simulator to complete its recorded route. When an agent reaches its final recorded position, it is removed from the simulator.
Short-term microscopic: Following [40, 5], we conduct a short-term microscopic evaluation by simulating for 20 seconds from a random time step in the test dataset. We first measure the similarity between the simulated and real data using position and velocity RMSE metrics given by:
@@
5.3 Performance
We compare our method against state-of-the-art baselines:
BC [32]: we learn our model directly by the BC method.
@@ 5.4 Ablation Study
We conducted a series of ablation experiments to assess the individual contributions of crucial components in our approach, whose results are presented in Tab. 1 and Tab. 2.
@@
On-road projection: Our ablation study on the on-road projection module aims to show its impact on reducing the off-road rate. The results show that the on-road projection module leads to a notable decrease in the off-road rate and moderate improvements in other performance metrics.
LQR: Removing LQR module allows us to evaluate its effectiveness. While LQR led to a higher short-term off-road rate due to more constrained movement (driving off-road due to inertia), its removal deteriorates other short-term and long-term metrics, because LQR can smooth the trajectory, thus leads to more realistic simulations.
@@ 5.5 Qualitative Result
In Fig. 2, we present the mean road density and speed for real-world data, SUMO simulation, and our proposed method over all time steps. The results demonstrate that our proposed method surpasses the capabilities of the SUMO simulator at replicating long-term macroscopic traffic patterns due to our model’s enhanced ability to replicate the typical microscopic driving behaviors over a long period.
@@ 6 Conclusion
7 Acknowledgements
This research project is partially supported by the Innovation and Technology Commission of the HKSAR Government under the InnoHK initiative, Hong Kong General Research Fund (11202119 and 11208718), Innovation and Technology Commission (GHP/126/21GD), and Guangdong, Hong Kong and Macao Joint Innovation Project (2023A0505010016).
@@ 8 Model details
The hyper-parameters of our model architecture are listed in Tab. 3. All models are trained using the Adam optimizer with a learning rate of 0.0003, and batch size 32 with 8 Ge-force 3090. The VAE are trained simultaneously with the policy network using the online simulated learned data and offline expert data.